# Are Users More Willing to Use Formally Verified Password Managers?

### v1

@@ 4.2. Results
Another aspect we explore is whether participants understand which features are formally verified. Specifically, we found that: (a) 60% identified the generator and the storage as formally verified; (b) 30% stated that the whole Password Manager was formally verified; when asked why one user stated “I saw the icon in several places” (P2); (c) and, one user (10%) could not explain, stating they did not know.
Users may be more willing to use a formally verified Password Manager.
When comparing users’ responses regarding trust in a formally verified Password Manager vs one that is not formally verified, users stated trusting more the formally verified Password Manager.
4.2.1. Limitations and Motivation for Second User Study.
A key limitation is the small sample size, which may not reflect a diverse population. Furthermore, our explanation of formal verification may have led participants to confuse security and verification. To avoid that issue in our main study, we completely refrain from explaining formal verification; instead, we communicate about the guarantees it provides. Another limitation is using a specific Password Manager, which might threaten the study’s generality. We argue that it did not influence the participants’ responses significantly as the interface does not appear remarkably different from other browser extension Password Managers (e.g., LastPass (lastpass, ) or 1Password (1password, )). However, we still decided not to use any specific Password Manager interface in our main study to facilitate generalization. After gathering the main insights from this explorative study, we designed and deployed a large-scale quantitative study informed by this formative study.
5. Main Study
To address all the research questions mentioned in Section 1 we designed and conducted a 200-participant online survey to understand users’ perceptions of formal verification.
Table 1. Percentage of participants that agreed or strongly agreed that the factor would impact their willingness to use a PM.
@@ 5.1. Structure
One of our goals was to understand if formal verification in a Password Manager affects users’ willingness to use it. To understand this, we first had to make sure that users knew what a Password Manager was, as previous studies have shown that some users do not use Password Managers because they do not know that these tools exist (pearman2019people, ; stobert2014password, ; alkaldi2016people, ). To explain what a Password Manager is and to situate the participants in our survey, we first present them with a vignette scenario where they read a news story about what a Password Manager is and the role of the primary password.
After viewing this scenario, we show participants a survey question addressing RQ1. We ask “Imagine that you were thinking about using a Password Manager. What would make you more willing to use the Password Manager?” We then present participants with different factors to which they respond using a Likert (lazar2017research, ) 5-point agreement scale. Among these factors, we include formal verification. The reason we added other factors in addition to formal verification was to: (1) Prevent participants from realizing that the focus of the study was formal verification. We did this to prevent common biases where the users are inclined to agree with the researcher (merrett2006reflections, ); (2) Understand if users’ preference for formally verified Password Managers differed significantly from other factors. These factors were, for example, that the Password Manager provides support materials such as tutorials and is free (see Table 1). One of the factors used was a certification by the “Password Manager Security Group”. This factor did not represent an actual entity. It was added to understand if participants were inclined to agree that any factor was impactful without understanding or because it sounded good.
Finally, we omit the term ”formal verification” by using analogies when possible. The specific analogy we chose is related to something most participants had to deal with in their education — mathematics. Instead of saying that the Password Manager is formally verified, we state that it is ”mathematically correct, that is, its features are as trustworthy as a mathematical proof.” We do this to follow best practices (redmiles2017summary, ) and exclude jargon from our survey as most users are unfamiliar with formal verification. Additionally, media outlets have previously used mathematical analogies to explain formal verification to broader audiences, such as the BBC (rubens_2015, ) and Quanta Magazine (hartnett_2020, ).
@@
After asking participants if they found that formal verification (among other factors) would impact their willingness to use a Password Manager, we asked them to elaborate. We presented participants with their answers and asked, respectively, ”In the previous question, you stated you felt more/less willing to use a Password Manager that is mathematically correct (…). Please state your reasons.”. We hoped to understand why participants valued (or not) formal verification in a Password Manager and what they associated it with. We also hoped to gauge if there was any misconception about what formal verification was for participants.
@@
To answer RQ2 and RQ3, we gather common Password Manager features (e.g., Password Generator and Clipboard clearing). For each of these, we present scenarios representing the impact formal verification can have on each feature. For example, for the Password Generator, the policy compliance scenario is: ”Imagine that you are creating a new account on a website (e.g., Twitter, Facebook). To increase security, you ask the Password Manager to generate a password with seven characters and at least two numbers. However, the password generated does not include any numbers.”. After each scenario, and using a 5-point Likert (lazar2017research, ) scale, we ask users if that scenario would make them stop using a Password Manager. A feature may have multiple scenarios if formal verification can impact it differently. An example is, again, the password generator, where formal verification may help with policy compliance but also with the guarantee that the generator is truly unpredictable (i.e., all passwords have the same probability). The scenarios were developed iteratively and followed guidelines for vignette scenarios design (barter1999use, ). In total, we designed seven unique scenarios.
5.2. Recruitment
@@
We recruited participants through Prolific. [^5] Prolific is a crowd-sourcing platform that enables large-scale user studies by connecting research and users https://prolific.com. We recruited 200 participants who were paid 2.65 GBP per submission (with one submission per user and an average time between 8 and 9 minutes). We deployed the with LimeSurvey. [^6] LimeSurvey is an open-source online survey tool https://www.limesurvey.org/.
Participant demographics
@@
About half of our sample identified as male and half as female. Our sample skewed younger than the general population as most participants were less than 34 years (87%), and the average age was 26. Most participants knew what a Password Manager was before the study (77.5%), but a slight majority of them stated that they did not use one currently (52%). The rest of the users were divided between those who said they were currently using a Password Manager (40%), those who had used one in the past but not anymore (3.5%), and those who did not know (3.5%). The three Password Managers that participants mentioned the most were Google Chrome Password Manager (41%), Bitwarden (15%), and Apple Keychain (14%).
@@
Nonetheless, due to a lack of understanding about what a Password Manager is some participants could answer that they had never used a Password Manager when they used a built-in Password Manager (e.g., when they received prompts to save a password in a browser). So, to understand the actual number of participants that use a Password Manager, we also asked participants if they saved their passwords in built-in Password Managers. As stated before, 52% of participants claimed they had never used a Password Manager before. However, of that 52%, 67% admitted to saving passwords when prompted (i.e., using built-in Password Managers). These results seem to imply that users use Password Managers without realizing they are using them. Reasons for these results could be a lack of technical knowledge, low computer literacy, and a lack of understanding about a Password Manager. We do not know if using a Password Manager without knowing impacts users’ perception of these tools. In total, 77% of participants used Password Managers to store and manage passwords.
5.3. Analysis
To understand if there was a significant difference between different survey questions, we used non-parametrical statistical tests as recommended by Lazar et al. (lazar2017research, ). We used Friedman’s ANOVA and Wilcoxon Signed-Rank Test with Bonferroni continuity correction (for repeated measures (lazar2017research, )). Non-nonparametric tests were used in previous studies on user perceptions of security issues (previous-usenix-nonparametrical, ).
@@ 5.3. Analysis
We coded the open answers from the survey using an inductive coding strategy. Two coders created the codebook iteratively using inductive and hierarchical coding (lazar2017research, ). In total, two coders coded 200 answers. We used Cohen’s Kappa to ensure inter-rater reliability and finished the coding process with a Cohen’s Kappa of 0.81, which denotes a very high level of agreement. An interpretation of Cohen’s Kappa is that a value above 0.60 indicates “satisfactory reliability” and one above 0.80 ”near-perfect agreement”  (lazar2017research, ). All answers were coded individually. However, regular meetings happened to discuss ambiguities and possible changes to the codebook. The final codebook consisted of 13 codes (see Appendix H), and each participant’s answer could have multiple codes. We used a hierarchical coding scheme for some of the codes, for example, the code “Extra Security” had as sub-codes: “Extra security: password generation”, for when participants mentioned that the Password Manager could create strong passwords; “Extra security: secure storage”, for when the participant stated that their passwords were safer inside the Password Managers’ storage; and ”Extra security: Protection from unwanted access”, for when they mentioned that the Password Manager prevents third-parties from accessing their passwords. We also included more general codes for when participants were not specific in their answers or showed a lack of understanding about what was being asked (”Answer general to all Password Managers and not specific to formally verified Password Managers” and ”Just seems better”) and one code for answers that were unusable to us (”Answer not relevant”). Two answers were coded as unusable, leaving this study with 198 usable participant answers for this question.
@@
Figure 3. Frequency of codes representing the reasons participants gave for ”agreeing” or ”disagreeing” that they felt willing to use a mathematically correct Password Manager.
5.4. Results
@@
5.4.1. Formal Verification and Willingness to use a Password Manager (RQ1)
@@
In the survey’s first question, we asked participants whether formal verification impacted their willingness to use a Password Manager. To prevent biases, we compared seven positive factors (see Table 1). The factors are all things that could be perceived as desirable to participants, so it is no surprise that most participants ”agreed” or ”strongly agreed” would be more likely to use a Password Manager with any of those factors. What we want to understand is if their opinion is the same across all factors or if there are some factors that users value more. A Friedman test showed that there was a statistically significant difference in participants’ answers between the different Password Manager’s features with a $\chi^{2}(2)=51.42$ and a $p<0.05$. However, the following pairwise comparison of users’ answers using the Wilcoxon rank sum testdid not find a significant difference between most factors ($p>0.05$).
@@
Comparing all the factors, we found that more than 86% of participants considered the use of formal verification (as described in Section 5.1) and being easy to use as important (they ”agreed” or ”strongly agreed” that these factors would make them more willing to use a Password Manager). The factors that users considered less critical (while still being essential for most) were being inexpensive (i.e., it costs money even if it is not expensive) and having support features (e.g., tutorials and help pages) with 31% and 27%, respectively, not thinking they would be more willing to use a Password Manager with them. This information can be found in Table 1. Our results suggest that users are more willing to use a formally verified Password Manager than a non-verified one. With this information in mind, we now aim to understand why participants answered our survey the way they did.
Understanding User’s Reasons.
This section analyzes the second survey section and aims to understand users’ reasons for varying formal verification. We separated the coding into the three questions that could be asked to participants — participants that valued formal verification, had no opinion, and found it undesirable (green, grey, and red in Figure 3). Additionally, depending on the size of the participant’s answer, one or more codes could be applied.
@@
Most participants who valued formal verification in their Password Managers mentioned the use of mathematics as a reason why (67 participants (35%)). Participants seem to associate mathematics with a “logical” and certain behavior. One even stated that they preferred to use a formally verified Password Manager “because mathematics doesn’t lie” (P89). Another stated, “I like to use such products or services that have scientific proof of their effectiveness” (P147).
@@
“Security” codes were also frequent (see Section 5.3). Some participants mentioned they believe formal verification added security (e.g., ”I think it would be more secure.” (P91)), and others gave more detailed explanations (e.g., “The main reason to start using a password manager is (…) the passwords generated are the strongest (…)” (P5)). Over 50% of the participants mentioned security as a reason why they valued formally verified Password Managers, with some mentioning more than one specific subcode of security in their answers (e.g., one participant stated that they preferred a formally verified Password Manager because ”…it will be more difficult to get hacked since it will create more secure passwords” (P162)).
@@
Overall, participants mostly agreed that formal verification was something to be desired in a Password Manager. They provided several reasons, but the more frequent ones were related to security and the use of mathematics. On the other hand, participants who did not find formal verification relevant had a general distrust of Password Managers, found that formal verification was not essential for them, or simply did not understand the concept. For example, one of the participants directly mentioned, ”I just don’t believe in that kind of app. ” (P22). Lack of understanding was also a frequent code present in the answers of participants who had a neutral response to this survey question (see the bars in yellow in Figure 3). Lack of understanding of technical concepts has been the cause of users’ trust issues with other security software in the past (e.g., in the domain of cryptocurrencies, lack of knowledge on the subject may cause users to question the security of these systems (presthus2017motivations, )). Nonetheless, most participants who agreed that formal verification was important to them stated that security had something to do with it.
@@
RQ1. How does formal verification impact users’ willingness to use Password Managers? Users seem to be more willing to use a formally verified Password Manager than a non-formally verified Password Manager; Our results also suggest that formal verification has positively affected some users’ trust; The main reasons users value formal verification are the use of mathematics and the extra security formal methods provide.
@@
Table 2. P-values of pairwise comparisons of users’ answers in the scenarios question using the Wilcoxon rank sum test and Bonferroni continuity correction. See Table 4 for the scenario descriptions.
@@
5.4.2. Formally Verified Features (RQ2)
After learning that formal verification seems to impact users’ interest in using Password Managers, we are interested in learning their priorities, thus addressing RQ2. We realize that users may not have the expertise to judge what is essential and what should be verified in a Password Manager. However, our goal is to understand what features, if verified, would be more impactful for users. We want to understand how to meet users’ wants and maximize the impact of formal verification. A Friedman’s test showed that there was a statistically significant difference in participants’ answers between the different scenarios with a $\chi^{2}(2)=331.08$ and a $p<0.05$. The results of a pairwise comparison of users’ answers using the Wilcoxon rank sum test and Bonferroni continuity correction can be seen in Table 2. Our results suggest that most of the differences between scenarios are statistically relevant ($p<0.05$).
Table 3. Percentage of participants that agreed or strongly agreed the scenario would make them stop using a PM. S# corresponds to Scenario # (see Table 4 in the Appendix for scenario descriptions).
@@ 5.4.2. Formally Verified Features (RQ2)
Overall, participants seem to consider some features more critical than others. Our results suggest that the secure vault and its encryption are the most important – 96% of participants stated they agreed or strongly agreed that they would stop using a Password Manager if the respective scenario (S3) occurred. In order the most important features to verify seem to be the secure vault (S3), password generator (S1 and S2), login in the Password Manager (i.e., primary password security in S4), clipboard (S6), synchronization across devices (S7 and S8) – see Table 3.
@@
RQ2. What features would users like to see formally verified in a Password Manager? The password vault’s security seems to be the feature that should be prioritized in a future implementation of a formally verified Password Manager; Surprisingly, verifying that the Password Manager’s passwords will not be lost does not seem to be important for users, as scenarios, where there was no third-party access to the password, were not as impactful; Ideally, all the features mentioned in this section should be formally verified, as over 50% of participants found their scenarios impactful.
5.4.3. Formal Verification Guarantees (RQ3)
The scenarios’ question also provides insights on whether participants value the guarantees of formal verification in Password Managers, thus answering RQ3.

### v2

@@ 4.2 Results
Another aspect we explore is whether participants understand which features are formally verified. Specifically, we found that: (a) 6 out of 10 identified the generator and the storage as formally verified; (b) 3 out of 10 stated that the whole PM was formally verified; when asked why one user stated “I saw the icon in several places” (P2); (c) and, one user could not explain, stating they did not know.
Users may be more willing to use a formally verified PM.
When comparing users’ responses regarding trust in a formally verified PM vs one that is not formally verified, users stated trusting more the formally verified PM.
4.2.1 Limitations and Motivation for Second User Study.
A key limitation is the small sample size, which may not reflect a diverse population. Furthermore, our explanation of formal verification may have led participants to confuse security and verification. To avoid that issue in our main study, we completely refrain from explaining formal verification; instead, we communicate about the guarantees it provides. Another limitation is using a specific PM, which might threaten the study’s generality. So, we decided not to use any specific PM interface in our main study. After gathering the main insights from this explorative study, we designed and deployed a large-scale quantitative study.
5 Main Study
To address all the RQs mentioned in Section˜2, we designed and conducted a 200-participant online survey to understand users’ perceptions of formal verification ( 8 in Figure˜1).
Table 1: Percentage of participants that agreed or strongly agreed that the factor would impact their willingness to use a PM.
@@ 5.1 Structure
One of our goals was to test whether formal verification in a PM affects users’ willingness to use it. First, we ensured participants knew what a PM was (prior work suggests users may not be familiar with PMs [41, 50].) To do this, we show participants a vignette scenario where they read a news story about what a PM is. After viewing this scenario, we show participants a survey question addressing RQ1. We ask “Imagine that you were thinking about using a PM. What would make you more willing to use the PM?” We then present participants with different factors to which they respond using a 5-point Likert scale  [33]. Among these factors, we include formal verification. The reason we added other factors in addition to formal verification was to prevent participants from realizing that the focus of the study was formal verification and prevent common biases where the users are inclined to agree with the researcher. We also wanted to understand if users’ preference for formally verified PMs differed from other factors (see Table˜1). One of the factors used was a certification by the “Password Manager Security Group”. This factor did not represent an actual entity. It was added to understand if participants were inclined to agree that any factor was impactful, without understanding, or because it sounded good.
Finally, we omit the term “formal verification” by using analogies when possible. The specific analogy we chose is related to something most participants had to deal with in their education — mathematics. Instead of saying that the PM is formally verified, we state that it is “mathematically correct, that is, its features are as trustworthy as a mathematical proof.” We do this to follow best practices [46] and exclude jargon from our survey, as most users are unfamiliar with formal verification. Media outlets have previously used mathematical analogies to explain formal verification to broader audiences (e.g., BBC [47] and Quanta Magazine [23]).
@@
After asking participants if they found that formal verification (among other factors), we presented participants with their answers and asked, respectively, “In the previous question, you stated you felt more/less willing to use a PM that is mathematically correct (…). Please state your reasons.” We hoped to understand why participants valued (or not) formal verification in a PM and what they associated it with.
@@
To answer RQ2 and RQ3, we gather common PM features (e.g., Password Generator and Clipboard clearing). For each of these, we present scenarios representing the impact formal verification can have on each feature. For example, for the Password Generator, the policy compliance scenario is: “Imagine that you are creating a new account on a website (e.g., Twitter, Facebook). To increase security, you ask the PM to generate a password with seven characters and at least two numbers. However, the password generated does not include any numbers.”. After each scenario, and using a 5-point Likert scale, we ask users if that scenario would make them stop using a PM. A feature may have multiple scenarios if formal verification can impact it differently. An example is, again, the password generator, where formal verification may help with policy compliance but also with the guarantee that the generator is truly unpredictable (i.e., all passwords have the same probability). We followed design guidelines [33] to iteratively develop seven unique scenarios
5.2 Recruitment
@@
We recruited participants through Prolific. [^2] Prolific is a crowd-sourcing platform that enables large-scale user studies by connecting research and users https://prolific.com. We recruited 200 participants who were paid 2.65 GBP per submission (with one submission per user and an average time between 8 and 9 minutes).
Participant demographics.
@@
About half of our sample identified as male and half as female. Our sample skewed younger than the general population, as most participants were less than 34 years (87%), and the average age was 26. Most participants knew what a PM was before the study (77.5%), but a slight majority of them stated that they did not use one currently (52%). The rest of the users were divided between those who said they were currently using a PM (40%), those who had used one in the past but not anymore (3.5%), and those who did not know (3.5%). The three PMs that participants mentioned the most were Google Chrome PM (41%), Bitwarden (15%), and Apple Keychain (14%).
@@
Nonetheless, due to a lack of understanding about what a PM is, some participants answered that they had never used a PM when they used a built-in PM (e.g., when they received prompts to save a password in a browser). We saw that, of the 52% of participants who had never used a PM, 67% of them saved passwords when prompted (e.g., in built-in PMs). These results seem to imply that users use PMs without realizing they are using them, maybe due to a lack of technical knowledge or low computer literacy. In total, 77% of participants used PMs to store and manage passwords.
5.3 Analysis
@@
To understand if there was a significant difference between different survey questions, we used Friedman’s test and Wilcoxon Signed-Rank Test with Bonferroni continuity correction [33]. We coded the open answers from the survey using an inductive coding strategy ( 12 in Figure˜1). Two coders created the codebook iteratively using inductive and hierarchical coding [33]. We coded 200 answers, achieving a Cohen’s Kappa of 0.81, indicating very high inter-rater reliability [33]. All answers were coded individually, and regular meetings happened to discuss ambiguities and updates to the codebook. The final codebook consisted of 13 codes and each participant’s answer could have multiple codes. We used a hierarchical coding scheme for some of the codes, for example, the code “Extra Security” had as sub-codes: “Extra security: password generation”, for when participants mentioned that the PM could create strong passwords and “Extra security: secure storage”, for when the participants stated that their passwords were safer inside the PMs’ storage. We also included more general codes for when participants were not specific in their answers or showed a lack of understanding about what was being asked (“Answer general to all PMs and not specific to formally verified PMs” and “Just seems better”).
@@
Figure 2: Frequency of codes representing the reasons participants gave for “agreeing” or “disagreeing” that they felt willing to use a mathematically correct PM.
5.4 Results
@@
5.4.1 Formal Verification and Willingness to use a PM (RQ1)
@@
In the survey’s first question, we asked participants whether formal verification impacted their willingness to use a PM. To reduce biases, we compared seven positive factors (see Table˜1). The factors are all things that could be perceived as desirable to participants, so it is no surprise that most participants “agreed” or “strongly agreed” would be more likely to use a PM with any of those factors. What we want to understand is if their opinion is the same across all factors or if there are some factors that users value more. A Friedman test showed that there was a statistically significant difference in participants’ answers between the different PM’s features with a $\chi^{2}(2)=51.42$ and a $p<0.05$. However, the following pairwise comparison of users’ answers using the Wilcoxon rank sum test did not find a significant difference between most factors ($p>0.05$).
@@
Comparing all the factors, we found that more than 86% of participants considered the use of formal verification (as described in Section˜5.1) and being easy to use as important (they “agreed” or “strongly agreed” that these factors would make them more willing to use a PM). The factors that users considered less critical (while still being essential for most) were being inexpensive (i.e., it costs money even if it is not expensive) and having support features (e.g., tutorials and help pages) with 31% and 27%, respectively, not thinking they would be more willing to use a PM with them. This information can be found in Table˜1. Our results suggest that users are more willing to use a formally verified PM than a non-verified one. With this information in mind, we now aim to understand why participants answered our survey the way they did.
Understanding Users’ Reasons.
@@
We separated answers into three categories: participants who valued formal verification, had no opinion, and found it undesirable (see Figure˜2). Most participants who valued formal verification in their PMs cited the use of mathematics as a reason (67 participants (35%)). Participants seem to associate mathematics with a “logical” and precise behavior. One even stated that they preferred to use a formally verified PM “because mathematics doesn’t lie” (P89). Another stated, “I like to use such products or services that have scientific proof of their effectiveness” (P147).
@@
“Security” codes were also frequent (see Section˜5.3). Some participants mentioned they believe formal verification added security (e.g., “I think it would be more secure.” (P91)), and others gave more detailed explanations (e.g., “The main reason to start using a PM is (…) the passwords generated are the strongest (…)” (P5)). Over 50% of the participants mentioned security as a reason why they valued formally verified PMs, with some mentioning more than one specific subcode of security in their answers (e.g., one participant stated that they preferred a formally verified PM because “… it will be more difficult to get hacked since it will create more secure passwords” (P162)).
@@
Overall, participants mostly agreed that formal verification was something to be desired in a PM. The most frequent reasons for this were related to security. On the other hand, participants who did not find formal verification relevant had a general distrust of PMs or found that formal verification was not essential for them. For example, one of the participants directly mentioned, “I just don’t believe in that kind of app” (P22). Lack of understanding was also a frequent code present in the answers of participants who had a neutral response to this survey question (see the bars in grey in Figure˜2). Lack of understanding of technical concepts has been the cause of users’ trust issues with other security software in the past [44]. Nonetheless, most participants who agreed that formal verification was important to them stated that security had something to do with it.
@@
RQ1. How does formal verification impact users’ willingness to use PMs? Users seem to be more willing to use a formally verified PM than a non-formally verified PM; Formal verification seems to have positively affected some users’ trust; Users primarily value formal verification for its reliance on mathematics and the additional security it provides.
5.4.2 Formally Verified Features (RQ2)
After learning that formal verification seems to impact users’ interest in using PMs, we are interested in learning their priorities, thus addressing RQ2. We realize that users may not have the expertise to judge what is essential and what should be verified in a PM. However, our goal is to understand what features, if verified, would be more impactful for users. We want to understand how to meet users’ wants and maximize the impact of formal verification. A Friedman’s test showed that there was a statistically significant difference in participants’ answers between the different scenarios with a $\chi^{2}(2)=331.08$ and a $p<0.05$. The results of a pairwise comparison of users’ answers using the Wilcoxon rank sum test and Bonferroni continuity correction can be seen in Table˜2. Our results suggest that most of the differences between scenarios are statistically relevant ($p<0.05$).
@@ 5.4.2 Formally Verified Features (RQ2)
Table 2: P-values of pairwise comparisons of users’ answers in the scenarios question using the Wilcoxon rank sum test and Bonferroni continuity correction.
@@
Table 3: Percentage of participants that agreed or strongly agreed the scenario would make them stop using a PM. S# corresponds to Scenario #
@@
Overall, participants seem to consider some features more critical than others. Our results suggest that the secure vault and its encryption are the most important — 96% of participants stated they agreed or strongly agreed that they would stop using a PM if the respective scenario (S3) occurred. The most important features to verify seem to be the secure vault (S3), password generator (S1), login in the PM (i.e., primary password security in S4), clipboard (S6), synchronization across devices (S7 and S8) (see Table˜3).
@@
RQ2. What features would users like to see formally verified in a PM? The password vault’s security seems to be the feature that should be prioritized in a future implementation of a formally verified PM; Surprisingly, verifying that the PM’s passwords will not be lost does not seem to be important for users, as scenarios where there was no third-party access to the password were not as impactful; Ideally, all the features mentioned in this section should be formally verified, as over 50% of participants found their scenarios impactful.
5.4.3 Formal Verification Guarantees (RQ3)